OpenStudy (anonymous):

A quality supervisor for a fctory that makes video games examines a sample of 82 video games. He finds that 7 video games are scratched. Which is the best prediction of the number of video games that will be scratched in a shipment of 6,000? A.500 B.512 C.556 D.610

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@Somy

OpenStudy (anonymous):

@Haseeb96

OpenStudy (anonymous):

@inowalst

OpenStudy (anonymous):

@chycora

OpenStudy (igreen):

Set up a proportion. \(\dfrac{7}{82} = \dfrac{x}{6000}\) Multiply 6000 to both sides: \(x = \dfrac{7 \times 6000}{82}\) Simplify

OpenStudy (anonymous):

Thank u :) @iGreen

OpenStudy (igreen):

No problem, what did you get?

OpenStudy (anonymous):

512? @iGreen

OpenStudy (igreen):

Correct!

OpenStudy (anonymous):

wow! yay! that you so much! @iGreen

OpenStudy (igreen):

No problem.